# DomainDemo: a dataset of domain-sharing activities among different demographic groups on Twitter

### v1

@@ Methods
We release two variants of the domain-sharing statistics dataset: DomainDemo-multivariate and DomainDemo-univariate. The table schema for DomainDemo-multivariate is provided in Table 1. For example, a row with the following values: domain=example.com, state=CA, race=Asian, gender=Female, age=30-49, party=Democrat, year_month=2018-12, shares=50, users=10, gini=0.1 indicates that the domain example.com was shared 50 times in December 2018 by 10 users who live in California, are Asian females aged between 30 and 49, and identify as Democrats. The Gini index of 0.1 indicates that the sharing count is almost evenly distributed across these users. DomainDemo-multivariate has 113,835,160 rows in total, with 129,127 unique domains. The dataset is divided into 132 parts, each corresponding to a month between May 2011 and April 2022. Note that party registration is not included in the multivariate variant.
Table 2: Schema of the DomainDemo-univariate tables. There are six universes: state, race, gender, age, party, and party registration, which share some common columns.
Data universe Column Data type Category Note
Shared by all universes domain string key second-level domains of the websites, e.g., nytimes.com
time string time year and month in the format “YYYY-MM”
shares integer statistic sharing event count of the domain in the bucket
users integer statistic unique number of users who share the domain in the bucket
state state string demographic two-letter aberrations, e.g., MA; including 50 U.S. states and the District of Columbia
race race string demographic can be one of “African-American,” “Caucasian,” “Hispanic,” “Asian,”“Other,” and “Unknown”
gender gender string demographic can be one of “Male,” “Female,” and “Unknown”
age age string demographic age bucket; can be one of “<18,” “18-29,” “30-49,” “50-64,” “65+,” and “Unknown”
party party string demographic inferred party; can be one of “Democrat,” “Independent,” and “Republican”
party registration partyreg string demographic party registration; can be one of “Democrat,” “Independent,” “Republican,” “Other,” and “Unknown”
DomainDemo-univariate details the statistics of sharing events across various categories within individual demographic variables. It includes separate sets of statistics (universes) for state, race, gender, age, party, and party registration. The table schema for these statistics is provided in Table 2. For example, a row in the state universe with the values: domain=example.com, year_month=2018-04, shares=3,000, users=250, gini=0.8, and state=NY indicates that the domain example.com was shared 3,000 times in April 2018 by 250 users in New York. The Gini index of 0.8 suggests that the sharing count is highly concentrated among a few users. Similar to DomainDemo-multivariate, each dataset within DomainDemo-univariate is divided into 132 monthly files.
@@
In addition to the domain-sharing statistics, we also release the derived metrics for domains. Note that the metrics can be computed at various temporal granularities, such as monthly intervals. For the released version, we calculate these metrics over the entire 11-year period. We provide the code to enable researchers to reproduce the metrics at different temporal granularities.
Domain localness
@@
We leverage the state universe data from DomainDemo-univariate to derive a data-driven metric that quantifies the “localness” of news domains. This is achieved by calculating the deviation of the user distribution of each domain in different states from the baseline distribution using Eq. (2). Details of the metric are provided in Table 3.
Domain audience partisanship
In this study, we employ DomainDemo-univariate and Eq. (3) to create data-driven metrics that assess the audience partisanship of domains by examining the party affiliation of the users who share these domains. It is important to note that our audience-based metrics do not evaluate the content characteristics of these domains. However, previous research indicates that audience characteristics are closely associated with the leanings of these domains (Gentzkow and Shapiro, 2010; Buntain et al., 2023).
We offer two versions of the audience partisanship scores: one based on inferred user partisanship and another based on party registration. Details of these metrics are provided in Table 3. The audience partisanship scores based on party registration cover fewer domains than other metrics due to the large number of missing values in the party registration column.
Other metrics
We also release three additional audience-based metrics: age deviation, race deviation, and gender leaning. These metrics can be used to understand the sharing patterns conditioned on these demographic variables. Details of the metrics are provided in Table 3.
Examples
Figure 1: Distributions of the derived metrics for all domains, along with detailed information for three domains cnn.com, news9.com, and wickedlocal.com. The left column presents the joint distributions of our derived metrics and the unique number of users who share each domain across the entire dataset. The color coding represents the number of domains within each grid cell. The symbols indicate the locations of the three domains. The three columns on the right provide detailed distributions of users in various demographic dimensions for the three domains, respectively. Sub-figures (b-d) highlight the discrepancies between the observed user distribution and the baseline distribution across U.S. states for the three domains. The color coding indicates the $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ value in each state. The bar plots display both the baseline distribution and the distribution of users in each demographic category for the domain of interest. The baseline distribution represents the patterns observed across all domains in the dataset.
To aid readers in comprehending the derived metrics, we present the distributions for all domains in the dataset and provide detailed information for three example domains in Figure 1. cnn.com, a national news outlet, has a user base closely aligned with the baseline. Consequently, its localness ($\mathcal{L}_{\delta}=0.013$), race deviation ($\mathcal{L}_{\delta}=0.002$), age deviation ($\mathcal{L}_{\delta}=0.049$), and gender leaning ($\mathcal{P}_{\delta}=-0.033$) scores are near zero. cnn.com is shared more often by Democratic users and less often by Republican users than the baseline, resulting in an audience partisanship score of $-0.132$.
Due to space constraints, we can only provide detailed demographic distributions for these three example domains. We have released an interactive app to allow readers to explore the patterns of other domains in our dataset at domaindemoexplorer.streamlit.app.
@@
In this section, we demonstrate the robustness of our dataset by first validating the demographic and partisanship measures from the voter records. We then compare the localness and audience partisanship scores of news domains against existing classifications.
Agreement with other data sources
In previous studies, our voter file measure of inferred partisanship was validated at the county level, and the race/ethnicity measure was compared against another voter file-based inference, yielding positive results (Shugars et al., 2021). Another investigation found that the demographic composition of our panel closely resembled representative survey samples of U.S. registered voters on Twitter (Hughes et al., 2021). At the individual level, the voter file measures were validated for 182 panel members also present in the survey samples, showing good matching for gender and reasonable matching for partisanship, but more significant mismatches for race/ethnicity minorities (Hughes et al., 2021). Overall, these findings demonstrate the general validity of the Twitter panel measures, and the domain-sharing statistics from our dataset can be considered generally representative of the sharing patterns of U.S. registered voters on Twitter.
However, the panel matching approach used in this work is not without risks. One source of error arises from the biases inherent in voter files. For instance, there are documented race/ethnicity disparities in who is listed in compiled voter records such as the TargetSmart data we use (Jackman and Spahn, 2021). Additionally, our race/ethnicity measure is inferred based on voter records for most states, implying a potential risk for biases in the categorization of individuals into race/ethnicity groups. Another potential source of error is the linkage process to Twitter users, where individuals with rare names are more likely to be included in the Twitter panel. In terms of representativity, the Twitter panel overrepresents white users and notably underrepresents Hispanics and Asians compared to the survey samples, and it also includes more females (Hughes et al., 2021). Therefore, our DomainDemo-univariate and DomainDemo-multivariate datasets will provide lower percentages of shares by racial minorities and men than a dataset fully representative of U.S. registered voters on Twitter would.
@@
Cronin et al. (2023) 1,390 276 1,666 Cronin et al. (2023) present a manually coded list of news outlets categorized as local, national, or international. We exclude international news sources from our study.
Fischer et al. (2020) 6,455 511 6,966 Fischer et al. (2020) offer a classification of news domains into categories such as local, regional, national, international, and technical. Their labels are derived from multiple sources, including Yin (2018) and manual annotation. For our analysis, we retain only the local and national domains. The authors do not provide specific criteria to distinguish regional from local categories and often combine them in their analysis. Therefore, we treat all regional domains as local in our study.
Horne et al. (2022) 284 0 284 Horne et al. (2022) compile a list of local news domains sourced from 50states.com/news.
Yin (2018) 6,493 0 6,493 Yin (2018) compile a list of local news outlets. The list is assembled from multiple sources, such as usnpl.com and stationindex.com.
@@
In addition to meta-ln, we compare our localness metric with that of Le Quéré et al. (2022). They also adopt a data-driven approach to quantify the localness of news domains from the audience perspective. Specifically, they quantify the “population reach” of news domains by measuring the distance between the locations of the outlets and the users following the outlets on Twitter while accounting for the population density. Since the population reach metric is continuous, we directly calculate its correlation with $\mathcal{L}_{\delta}$. The intersection between the list shared by Le Quéré et al. (2022) and ours has 1,342 domains and yields a Spearman correlation coefficient of 0.441 ($p<0.001$), suggesting a moderate agreement between the two metrics.
Although our focus here is on news outlets, our dataset encompasses a much wider variety of websites. For instance, our dataset includes information on news domains lacking clear editorial guidelines, such as misinformation websites and “pink slime” websites (Moore et al., 2023). It also covers non-news websites, including government, organizational, entertainment, and e-commerce sites.
@@
Eady et al. (2020) Media ideology scores based on Twitter data. The authors jointly estimate the idealogy of politicians, users and news sources through the news sharing behaviors on Twitter. 179 0.929 0.929
@@
Allsides (allsides.com/media-bias/media-bias-ratings) Allsides produces domain bias scores based on their own algorithm. We map their labels “left,” “left-lean,” “center,” “right-lean,” and “right” to numerical values $-1$, $-0.5$, 00, $0.5$, $1$ for analysis. 189 0.736 0.743
@@
Data and code availability
Given the sensitive nature of the information about Twitter users in our datasets, we have implemented layered access controls. Since DomainDemo-multivariate can potentially reveal the identities of Twitter users in the dataset when combined with other datasets, it will be hosted within a secure platform. To access this dataset, researchers must complete an application process and sign a data use agreement that prohibits the identification of individual Twitter users. The highly aggregated nature of DomainDemo-univariate reduces the risk of re-identification. Thus, after completing an application process, researchers will be able to download these data files. The derived metrics of the domains, such as localness and audience partisanship scores, are made publicly available.
The details on how to access DomainDemo-multivariate and DomainDemo-univariate will be released soon. At this time, we have released the derived metrics of the domains, together with the code to compute them at github.com/LazerLab/DomainDemo. We also provide an app that allows exploring the derived metrics interactively at domaindemoexplorer.streamlit.app.
Acknowledgements
The authors thank Lisa Friedland for assisting with the data collection and Ronald Robertson for providing the validation data.

### v2

@@ Methods
domains_count_mean float statistic average number of unique domains shared by users in the bucket
domains_count_std float statistic standard deviation of the number of unique domains shared by users in the bucket
state string demographic two-letter abbreviations, e.g., MA; including 50 U.S. states and the District of Columbia; only included in the “state” universe
race string demographic can be one of “African-American,” “Caucasian,” “Hispanic,” “Asian,”“Other,” and “Unknown”; only included in the “race” universe
gender string demographic can be one of “Male,” “Female,” and “Unknown”; only included in the “gender” universe
age string demographic age bucket; can be one of “<18,” “18-29,” “30-49,” “50-64,” “65+,” and “Unknown”; only included in the “age” universe
party string demographic inferred party; can be one of “Democrat,” “Independent,” and “Republican”; only included in the “party” universe
partyreg string demographic party registration; can be one of “Democrat,” “Independent,” “Republican,” “Other,” and “Unknown”; only included in the “party registration” universe
DomainDemo-univariate details the statistics of sharing events across various categories within individual demographic variables. It includes separate sets of statistics (universes) for state, race, gender, age, party, and party registration. The table schema for these statistics is provided in Table 2. A row in the state universe monthly distribution variant with the values: domain=example.com, year_month=2018-04, shares=3,000, users=250, gini=0.8, and state=NY indicates that the domain example.com was shared 3,000 times in April 2018 by 250 users in New York. The Gini index of 0.8 suggests that the sharing count is highly concentrated among a few users. A row in the baseline variant with the values: state=NY, year_month=2018-04, shares=451,000, users=18,250, gini=0.5, domains_count_mean=10.9, domains_count_std=10.2 indicates that there were 451,000 sharing events of any domains in April 2018 by 18,250 users in New York. These users shared an average of 10.9 unique domains with a standard deviation of 10.2.
@@
In addition to the domain-sharing statistics, we also release the derived metrics for domains based on the data from the whole time period. Each of our released files contains two columns: domain and the corresponding metric value. Details of these metrics are provided in Table 3. Note that we offer two versions of the audience partisanship scores: one based on inferred user partisanship and one based on party registration. The audience partisanship scores based on party registration cover fewer domains than other metrics due to the missing values of the party registration information.
@@
In this section, we discuss the robustness of the demographic variables in our dataset. We then compare the localness and audience partisanship scores of news domains against existing classifications.
Demographic Variables
The demographic variables of Twitter panel users are the foundation of DomainDemo. While other information is self-reported, the partisanship score and race are inferred by TargetSmart. Although the inference algorithms remain proprietary, multiple lines of evidence support their reliability. For the partisanship score, we find that it highly correlates with the party registration information for individuals registered as Democrats or Republicans in the 30 states plus the District of Columbia where party registration information is available, with a 94% agreement rate. An independent evaluation from the Pew Research Center also suggests this inferred partisanship is reasonably accurate (Igielnik et al., 2018). Moreover, our previous research validates these scores through their strong alignment with county-level election results (Shugars et al., 2021). Similarly, TargetSmart’s race estimates show consistency with different reference points, including self-reported race data from a Pew Research Center survey and results from a statistical inference method (Shugars et al., 2021).
The representativeness of the Twitter panel is another important aspect of our dataset. Our previous research has compared the panel with a representative sample of registered voters on Twitter created by the Pew Research Center (Hughes et al., 2021). The study shows substantial agreement between the two samples in general, despite some noteworthy differences. In particular, the panel exhibits an overrepresentation of Caucasian users while underrepresenting other racial groups, particularly Hispanic and Asian populations. Additionally, the panel contains a slightly higher share of female users and younger individuals compared to the survey samples.
Figure 3: Comparison of the demographic composition of the Twitter panel with that of all registered voters. The age is calculated using 2017 as the reference year.
Here, we further compare the demographic composition of the Twitter panel with all registered voters in the TargetSmart voter file, as illustrated in Figure 3. Notably, Twitter panel users tend to be younger than registered voters. For other aspects, the panel generally reflects the composition of registered voters with some minor differences. In particular, the panel contains a higher proportion of Caucasian users while underrepresenting other racial groups. Additionally, male users and Republican users are slightly underrepresented in the Twitter panel.
These comparative analyses offer valuable insights into the representativeness of our Twitter panel. The panel demonstrates reasonable alignment with the broader population of registered voters and voters on Twitter. However, researchers should exercise caution when interpreting results, particularly regarding potential biases in age distribution, gender representation, and especially racial composition.
@@
Cronin et al. (2023) 1,390 276 1,666 Cronin et al. present a manually coded list of news outlets categorized as local, national, or international. We exclude international news sources from our study.
Fischer et al. (2020) 6,455 511 6,966 Fischer et al. offer a classification of news domains into categories such as local, regional, national, international, and technical. Their labels are derived from multiple sources, including Yin et al. and manual annotation. For our analysis, we retain only the local and national domains. The authors do not provide specific criteria to distinguish regional from local categories and often combine them in their analysis. Therefore, we treat all regional domains as local in our study.
Horne et al. (2022) 284 0 284 Horne et al. compile a list of local news domains sourced from 50states.com/news.
Yin (2018) 6,493 0 6,493 Yin et al. compile a list of local news outlets. The list is assembled from multiple sources, such as usnpl.com and stationindex.com.
@@
In addition to meta-ln, we compare our localness metric with that of Le Quéré et al. (2022). Le Quéré et al. also adopt a data-driven approach to quantify the localness of news domains from the audience perspective. Specifically, they quantify the “population reach” of news domains by measuring the distance between the locations of the outlets and the users following the outlets on Twitter while accounting for the population density. Since the population reach metric is continuous, we directly calculate its correlation with $\mathcal{L}_{\delta}$. The intersection between the list shared by Le Quéré et al. and ours has 1,342 domains and yields a Spearman correlation coefficient of 0.441 ($p<0.001$), suggesting a moderate agreement between the two metrics.
@@
Eady et al. (2025) Media ideology scores based on Twitter data. The authors jointly estimate the ideology of politicians, users, and news sources through the news sharing behaviors on Twitter. 179 0.929 0.929
@@
Allsides (allsides.com/media-bias/media-bias-ratings) Allsides produces domain bias scores based on its own algorithm. We map their labels “left,” “left-lean,” “center,” “right-lean,” and “right” to numerical values $-1$, $-0.5$, 00, $0.5$, $1$ for analysis. 189 0.736 0.743
@@
Usage Notes
Our dataset offers a comprehensive view of domain sharing patterns on Twitter, capturing variations across demographic groups throughout an extended period. The demographic characteristics of the audiences also reveal distinctive patterns that illuminate the nature of the shared domains.
A key application of our dataset lies in examining the U.S. news media landscape. For researchers interested in this area, we refer them to a curated list of news domains (Yang, 2025). Researchers can integrate this list with our dataset to identify and analyze news domains within our collection.
Beyond news media, our dataset encompasses a diverse range of web domains. This includes news-like websites without established editorial standards, such as misinformation sites and “pink slime” websites (Moore et al., 2023). The dataset also extends to various non-news domains, spanning government websites, organizational platforms, entertainment sites, and e-commerce portals.
Code Availability
The code associated with this dataset is available at https://github.com/LazerLab/DomainDemo. We share example scripts to load and analyze the data in DomainDemo. We also provide code to reproduce the derived metrics for domains and the validation results.